# Shoot from the HIP: Hessian Interatomic Potentials Without Derivatives

### v1

@@ 3.3 Loss function design
However, we are often mainly interested in the smallest eigenvalues and the corresponding eigenvectors (Besalú & Bofill, 1998; Cerjan & Miller, 1981). For this reason, we design a loss function to emphasize this subspace of the eigen-decomposition.
Let $\mathbf{V}=[\mathbf{v}_{1},\mathbf{v}_{2},...,\mathbf{v}_{3N}]$ be the matrix with columns made from eigenvectors of the ground truth Hessian $\mathbf{H}$, and $\mathbf{\Lambda}$ the corresponding matrix with eigenvalues on the diagonal and zeros everywhere else. Further denote $\mathbf{V}_{[:,:k]}$ and $\mathbf{V}_{[:,k:]}$ the matrices sliced to contain all columns up to/starting from the $k^{th}$ one. We can then define a subspace loss by
@@
If $\mathbf{H}^{\text{pred}}=\mathbf{H}$ we have $\mathbf{V}^{\top}\mathbf{H}^{\text{pred}}\mathbf{V}=\mathbf{\Lambda}$, and the loss has therefore the correct minimum. Using $\mathcal{L}=\mathcal{L}_{\text{MAE/MSE}}+\alpha\mathcal{L}_{\text{sub}}$ lets us emphasize the subspace corresponding to the $k$ lowest eigenvectors and eigenvalues. Since the Hessian has six redundant degrees of freedom with eigenvalues close to zero, we need to set $k>6$. In practice, we use $k=8$. As we will show in A.5, predicting Hessians with HIP using any loss function significantly outperforms AD Hessians. MAE combined with the subspace loss in equation 10 further improves results in transition state search and extrema classification with frequency analysis by a small margin, which is why we use them in the experiment of the main text. For details and ablations of the choice of loss function see A.5.
@@ 4 Experiments
Since DFT Hessians are computationally expensive to calculate for even small molecules, there are not many available datasets. Most notable are the Hessian-QM9 (Williams et al., 2025) and the HORM (Cui et al., 2025) datasets, with the latter being used in this study due to its larger proportion of non-equilibrium molecules compared to Hessian-QM9.
@@
To quantitatively assess the accuracy and speed of our directly predicted Hessians, we calculate the average elementwise mean absolute error, average eigenvalue error, and average time per prediction. Since, for many tasks such as ZPE and TS search, the smallest eigenvalue/eigenvector pair is particularly important, we calculate the cosine similarity and eigenvalue error for these tasks separately. The results are presented in Table 1. In every metric, our direct prediction Hessian approach outperforms the other models. The improvement is especially notable in prediction speed, where we are an order of magnitude faster than the models relying on automatic differentiation.
In Figures 2(a) and (b), we visually inspect the HIP Hessian compared to the ground-truth DFT Hessian, respectively. We observe good agreement with the absolute difference between the two shown in Figure 2(c).
Table 1: Comparison of Hessian prediction errors on the HORM-Transition1x validation set. Numbers denote the mean absolute error (MAE), except for the cosine similarity (CosSim) and average time per forward pass of a single molecule. Bold values highlight the best-performing model in each task. Baseline AD models are taken from (Cui et al., 2025).
@@
AD AlphaNet 0.385 2.871 0.750 0.770 767.0
LEFTNet 0.150 0.684 0.804 0.322 1110.7
LEFTNet-df 0.197 0.669 0.532 0.742 341.3
EquiformerV2 0.074 0.242 0.541 0.324 633.0
Predicted HIP-EquiformerV2 0.030 0.063 0.870 0.130 38.5
Figure 2: HIP compared to ground-truth DFT Hessian. The horizontal and vertical axis index the $(3N_{Atoms})^{2}$ entries of the Hessian. We depict the absolute values of test sample 1234512345 in eV/Å2.
@@
Figure 3: Computational cost. (a) Speed and (b) memory footprint for a single sample as a function of molecule size, and (c) scaling of Hessian predictions in parallel with increasing batch size. AD values are shown in blue, and HIP (this work) in orange.
In Figure 3(a-b), we further compare the prediction time and memory footprint as a function of molecule size. The direct prediction of Hessians is $10$ - $74\times$ faster than using AD for molecules in the HORM dataset. HIP predicted Hessians also benefit from a much more favorable scaling with the molecule size. Figure 3(c) compares the batched AD Hessian as implemented in (Cui et al., 2025) with our HIP-EquiformerV2. We observe at minimum a $78\times$ speedup as a function of batch size compared to AD. We explain the significant performance degradation of batched AD Hessians in A.3.
@@
We now turn to downstream applications of the Hessian, first focusing on geometry optimization. We compare exact second-order approaches using Hessians (RFO) against first-order methods, and quasi-second-order RFO initialized with Hessians followed by BFGS updates. See A.6.1 for details. Figure 4(a) shows that RFO with HIP Hessians converges within the fewest steps in the median case. In contrast, both finite difference Hessians and AD Hessians frequently fail to converge. Among the Hessian-initialized BFGS methods (blue), the predicted Hessian also performs the best. Steepest descent with line search does not converge within the step budget, which demonstrates the difficulty of the task.
As seen in Figure 4(b), RFO and BFGS methods with HIP Hessians are also the fastest in terms of wall-clock time. The current implementation is bottlenecked by memory transfer from and to the CPU. We would expect an even more significant speedup if the optimization was performed on GPU and parallelized across samples.
@@
We evaluated an important thermochemical property of the relaxed equilibrium state, the zero-point energy (ZPE). We report both the ZPE of the reactant, as well as the relative ZPE between reactant and product $\Delta\mathrm{ZPE}$. Experimental details are in A.6.3. Table 2 shows the mean absolute error with standard deviation (MAE (Std)) for each model compared to DFT. The HIP-predicted Hessians reach an order of magnitude lower MAE than the next-best model, and two orders of magnitude lower MAE than the same model using automatic differentiation Hessians. The results are also well below the chemical accuracy of 43 meV.
Transition state search
Another important application of Hessians is transition state search. We use the recently introduced ReactBench benchmark (Zhao et al., 2025), but additionally verify the predicted transition states with DFT. A step-by-step description of the workflow can be found in A.6.2. We present metrics of each step of the workflow in Figure 6. The growing string method (GSM) success rate is the same for both AD and HIP-EquiformerV2 models. GSM only uses energy and forces, which are the same for both models. For all subsequent metrics, the HIP predicted Hessians outperform the AD ones. The RFO convergence and DFT-verified convergence metrics depend on the Hessian the most. For both, the HIP Hessians yield the most improvements over AD Hessians.
@@
Direct Hessian regression requires ground-truth data from DFT, which can be particularly expensive to obtain for larger systems. Therefore, the experiments in this work focus on small organic molecules. Future work should explore applying Hessian prediction to other chemical systems, such as materials, larger complexes, or even proteins. Furthermore, to isolate the Hessian from the force and energy predictions, we trained only the Hessian regression head using a fixed backbone. For optimal results, energy, forces, and Hessians should be trained end-to-end.
@@ A.1.3 Geometry optimization
Geometry optimization locates minima $\operatorname*{arg\,min}_{\mathbf{R}}E(\mathbf{R})$ of the potential energy surface starting from non-equilibrium molecular geometries. It underpins most workflows from global searches to high-throughput screening. The efficiency of geometry optimization depends on four key factors: (i) the choice of coordinate system, (ii) the quality of the Hessian approximation, (iii) the update strategy, and (iv) the step size control. Cartesian coordinates are straightforward but strongly coupled, which limits step size and slows convergence for flexible systems. In contrast, internal coordinates are better aligned with chemically meaningful motions, separate stiff and soft modes, and allow constraints to be imposed naturally. These properties often lead to substantially fewer iterations. Some of the most widely used optimizers are the first-order FIRE (Bitzek et al., 2006) optimizer and the second-order rational function optimization (RFO) optimizer.
@@
Rational function optimization (RFO) (Simons et al., 1983; Banerjee et al., 1985) is a commonly used second-order optimization technique in molecular geometry optimization. RFO starts with a [2/2]-Padé-expansion of the energy:
@@
An attractive property of RFO is that it can be used both for minimizing $E(\mathbf{R})$ as well as for finding saddle points (important for transition state search, see next section): If we pick the first eigenvector of equation 20, we get an update pointing to the minimum; if we select the second eigenvector, we get a transition state update. In contrast to Newton-Raphson optimization, RFO is also robust to indefinite Hessians. For detailed derivation, see (Simons et al., 1983; Banerjee et al., 1985). In this paper, for minimization, we are using restricted step RFO (RS-RFO), which adds a trust region to prevent unphysically large steps. For the transition point search, we are using restricted step partitioned RFO (RS-P-RFO), a slight modification that treats the subspace with negative and positive eigenvalues of the Hessian separately.
@@ A.1.4 Transition State Search
Computational methods for transition state search can broadly be categorized as single-ended and double-ended. In double-ended methods, we know the product and reactant states (the two minima on the energy surface) and try to find an interpolation on the MEP that goes through the TS. In contrast, in single-ended methods, we only know a starting point and try to climb up on the energy surface to find a nearby transition state. Often, double-ended methods are used to give good initial guesses, which are then refined by single-ended methods. In this study, we are using the Growing string method to find the initial states, which we then refine with RS-P-RFO (see above).
Different double-ended methods are distinguished by how they grow the interpolations between product and reactant states. The most common approaches are the nudge elastic band method (Henkelman & Jónsson, 1999; Henkelman et al., 2000) and the growing string method (GSM) (Peters et al., 2004; Zimmerman, 2015).
@@
The nudged elastic band method aims to find the minimum energy path between given initial and final configurations. A set of images connecting these states is linked by classical spring forces, forming an elastic band. Each image experiences a total force comprising the spring force along the local tangent and the true force perpendicular to it. The images are simultaneously optimized to trace the MEP. To converge to the transition state, a climbing image is introduced, for which the force along the band is inverted, leaving only the perpendicular component of the true force. This drives the climbing image up the potential energy surface along the band while descending perpendicularly, reaching the transition state along the MEP.
@@
where $\mathbf{F}=-\nabla E$ is the force and $\hat{\mathbf{t}}$ is the local tangent unit vector. Convergence is typically accelerated by approximate Hessians, constructed in delocalized internal coordinates and updated by quasi-Newton schemes, which enable eigenvector-guided optimization while avoiding full second-derivative evaluations (Zimmerman, 2015). Once the two fragments merge, the full string is reparameterized to maintain uniform node distance, and the highest energy node serves as a transition state estimate, which can be further refined using Hessian-based eigenvector following (Peters et al., 2004).

### v2

@@ 3.3 Loss function design
However, we are often mainly interested in the smallest eigenvalues and the corresponding eigenvectors (RSRFO1998a; cerjanFindingTransitionStates1981). For this reason, we use a loss function that emphasize the relevant subspace of Hessian. This loss function is similar to li2025enhancing, which uses it to emphasize the subspace up to the Lowest Unoccupied Molecular Orbital (LUMO) in Fock matrix prediction. Let $\mathbf{V}=[\mathbf{v}_{1},\mathbf{v}_{2},...,\mathbf{v}_{3N}]$ be the matrix with columns made from eigenvectors of the ground truth Hessian $\mathbf{H}$, and $\mathbf{\Lambda}$ the corresponding matrix with eigenvalues on the diagonal and zeros everywhere else. Further denote $\mathbf{V}_{[:,:k]}$ and $\mathbf{V}_{[:,k:]}$ the matrices sliced to contain all columns up to/starting from the $k^{th}$ one. We can then define a subspace loss by
@@
If $\mathbf{H}^{\text{pred}}=\mathbf{H}$ we have $\mathbf{V}^{\top}\mathbf{H}^{\text{pred}}\mathbf{V}=\mathbf{\Lambda}$, and the loss has therefore the correct minimum. Using $\mathcal{L}=\mathcal{L}_{\text{MAE/MSE}}+\alpha\mathcal{L}_{\text{sub}}$ lets us emphasize the subspace corresponding to the $k$ lowest eigenvectors and eigenvalues. Since the Hessian always has six redundant degrees of freedom with eigenvalues close to zero, we need to set $k>6$. In practice, we use $k=8$. As we will show in A.5.1, predicting Hessians with HIP using any loss function significantly outperforms AD Hessians. MAE combined with the subspace loss in equation 9 further improves results in transition state search and extrema classification with frequency analysis by a small margin, which is why we use them in the experiment of the main text. For details and ablations of the choice of loss function see A.5.1.
@@ 4 Experiments
To quantitatively assess the accuracy and speed of our directly predicted Hessians, we calculate the average elementwise mean absolute error, average eigenvalue error, and average time per prediction. Since, for many tasks such as ZPE and TS search, the smallest eigenvalue/eigenvector pair is particularly important, we calculate the cosine similarity and eigenvalue error for these tasks separately. The results are presented in Table 1. Models trained on just energy and forces without Hessian information fail completely (EquiformerV2 (E-F) and cui2025horm). In every metric, our direct prediction Hessian approach outperforms the other models. The improvement is especially notable in prediction speed, where we are an order of magnitude faster than the models relying on automatic differentiation. Training HIP end-to-end improves the results further.
@@
AD EquiformerV2 (E-F) 2.252 20.752 0.279 1.372 613.3
AlphaNet 0.385 2.871 0.900 0.284 767.0
LEFTNet-CF 0.150 0.684 0.943 0.112 1110.7
LEFTNet-DF 0.197 0.669 0.930 0.148 341.3
EquiformerV2 0.074 0.242 0.915 0.097 633.0
Predicted HIP-EquiformerV2 0.030 0.063 0.982 0.031 38.5
HIP-EquiformerV2 (end-to-end) 0.020 0.041 0.982 0.031 31.4
Table 1: Comparison of Hessian prediction errors on the HORM-T1x validation set. Denoted are the mean absolute errors (MAE), cosine similarity (CosSim), and average time per forward pass on a RTX 3060 GPU. $\lambda_{1}$ and ${\bm{v}}_{1}$ were obtained after Eckart-projection.
@@
Figure 2: Computational cost. (a) Speed and (b) memory footprint for a single sample as a function of molecule size, and (c) scaling of Hessian predictions in parallel with increasing batch size. Timings were performed on an RTX 3060 GPU. AD values are shown in blue, and HIP (this work) in orange.
In Figure 2(a-b), we further compare the prediction time and memory footprint as a function of molecule size. The direct prediction of Hessians is $10$ - $74\times$ faster than using AD for molecules in the HORM dataset. HIP predicted Hessians also benefit from a much more favorable scaling with the molecule size. Figure 2(c) compares the batched AD Hessian as implemented in (cui2025horm) with our HIP-EquiformerV2. We observe at minimum a $78\times$ speedup as a function of batch size compared to AD. We explain the significant performance degradation of batched AD Hessians in A.3.
@@
We now turn to downstream applications of the Hessian, first focusing on geometry optimization. We compare exact second-order approaches using Hessians (RFO) against first-order methods, and quasi-second-order RFO initialized with Hessians followed by BFGS updates. See A.6.1 for details. Figure 3(a) shows that RFO with HIP Hessians converges within the fewest steps in the median case. In contrast, both finite difference Hessians and AD Hessians frequently fail to converge. Among the Hessian-initialized BFGS methods (blue), the predicted Hessian also performs the best. Steepest descent with line search does not converge within the step budget, which demonstrates the difficulty of the task. As seen in Figure 3(b), RFO and BFGS methods with HIP Hessians are also the fastest in terms of wall-clock time.
@@
We evaluated an important thermochemical property of the relaxed equilibrium state, the zero-point energy (ZPE). We report both the ZPE of the reactant, as well as the relative ZPE between reactant and product $\Delta\mathrm{ZPE}$. Experimental details are in A.6.3. Table 2 shows that the HIP-predicted Hessians reach an order of magnitude lower MAE than the next-best model, and two orders of magnitude lower MAE than the same model using automatic differentiation Hessians.
@@
Transition state search
Another important application of Hessians is transition state search. We use the recently introduced ReactBench benchmark (zhao2025), but additionally verify the predicted transition states with DFT. A step-by-step description of the workflow can be found in A.6.2. We present metrics of each step of the workflow in Figure 5. The growing string method (GSM) success rate is the same for both AD and HIP-EquiformerV2 models. GSM only uses energy and forces, which are the same for both models. For all subsequent metrics, the HIP predicted Hessians outperform the AD ones. The RFO convergence and DFT-verified convergence metrics depend on the Hessian the most. For both, the HIP Hessians yield the most improvements over AD Hessians.
@@
End-to-end HIP training improves energies and forces
So far we restricted ourselves to only finetuning the Hessian prediction head while freezing the existing parameters. We now investigate the more realistic decision of training either a regular MLIP on energy and forces, or a HIP including the Hessian prediction. We train both model types for the same wallclock time of 3 days at various data scales. Figure 6 shows that end-to-end training of the joint backbone and the readout not only improves the Hessian, but also the energy and force predictions.
Figure 6: Accuracy scaling with dataset size. (a) Energy (b) Force (c) Hessian MAE on the HORM-T1x validation over the number of training samples randomly chosen from the HORM-T1x train set. EquiformerV2 trained on energy and forces is shown in dark gray, the same model trained as HIP also on Hessian data (this work) in orange.
@@
Direct Hessian regression requires ground-truth data from DFT, similar to the energy and force labels used for regular MLIPs. The experiments in this work focus on small organic molecules, and future work should explore applying Hessian prediction to other chemical systems, such as materials, larger complexes, or even proteins. To isolate the Hessian from the force and energy predictions, we trained only the Hessian regression head using a fixed backbone. For optimal results on all benchmarks, energy, forces, and Hessians should be trained end-to-end.
@@ A.1.3 Geometry optimization
Geometry optimization locates minima $\operatorname*{arg\,min}_{\mathbf{R}}E(\mathbf{R})$ of the potential energy surface starting from non-equilibrium molecular geometries. It underpins most workflows from global searches to high-throughput screening. The efficiency of geometry optimization depends on four key factors: (i) the choice of coordinate system, (ii) the quality of the Hessian approximation, (iii) the update strategy, and (iv) the step size control. Cartesian coordinates are straightforward but strongly coupled, which limits step size and slows convergence for flexible systems. In contrast, internal coordinates are better aligned with chemically meaningful motions, separate stiff and soft modes, and allow constraints to be imposed naturally. These properties often lead to substantially fewer iterations. Some of the most widely used optimizers are the first-order FIRE (bitzek2006) optimizer and the second-order rational function optimization (RFO) optimizer.
@@
Rational function optimization (RFO) (simons1983; banerjee1985) is a commonly used second-order optimization technique in molecular geometry optimization. RFO starts with a [2/2]-Padé-expansion of the energy:
@@
An attractive property of RFO is that it can be used both for minimizing $E(\mathbf{R})$ as well as for finding saddle points (important for transition state search, see next section): If we pick the first eigenvector of equation 19, we get an update pointing to the minimum; if we select the second eigenvector, we get a transition state update. In contrast to Newton-Raphson optimization, RFO is also robust to indefinite Hessians. For detailed derivation, see (simons1983; banerjee1985). In this paper, for minimization, we are using restricted step RFO (RS-RFO), which adds a trust region to prevent unphysically large steps. For the transition point search, we are using restricted step partitioned RFO (RS-P-RFO), a slight modification that treats the subspace with negative and positive eigenvalues of the Hessian separately.
@@ A.1.4 Transition State Search
Computational methods for transition state search can broadly be categorized as single-ended and double-ended. In double-ended methods, we know the product and reactant states (the two minima on the energy surface) and try to find an interpolation on the minimum energy path (MEP) that goes through the TS. In contrast, in single-ended methods, we only know a starting point and try to climb up on the energy surface to find a nearby transition state. Often, double-ended methods are used to give good initial guesses, which are then refined by single-ended methods. In this study, we are using the Growing string method to find the initial states, which we then refine with RS-P-RFO (see above).
Different double-ended methods are distinguished by how they grow the interpolations between product and reactant states. The most common approaches are the nudge elastic band method (henkelman1999; henkelman2000) and the growing string method (GSM) (peters2004; zimmerman2015).
@@
The nudged elastic band method aims to find the minimum energy path between given initial and final configurations. A set of images connecting these states is linked by classical spring forces, forming an elastic band. Each image experiences a total force comprising the spring force along the local tangent and the true force perpendicular to it. The images are simultaneously optimized to trace the minimum energy path (MEP). To converge to the transition state, a climbing image is introduced, for which the force along the band is inverted, leaving only the perpendicular component of the true force. This drives the climbing image up the potential energy surface along the band while descending perpendicularly, reaching the transition state along the MEP.
@@
where $\mathbf{F}=-\nabla E$ is the force and $\hat{\mathbf{t}}$ is the local tangent unit vector. Convergence is typically accelerated by approximate Hessians, constructed in delocalized internal coordinates and updated by quasi-Newton schemes, which enable eigenvector-guided optimization while avoiding full second-derivative evaluations (zimmerman2015). Once the two fragments merge, the full string is reparameterized to maintain uniform node distance, and the highest energy node serves as a transition state estimate, which can be further refined using Hessian-based eigenvector following (peters2004).
@@ A.2 Hessian properties
Non-integrability Hessians
Since direct Hessian predictions are not formulated as derivatives, they are not guaranteed to be integrable. It is important to understand when a non-integrability error can lead to catastrophically wrong downstream predictions similar to direct force predictions in MD simulations.
Lets first review the more familiar case: Why do direct force predictions fail in MD simulations but succeed in geometry optimization? Suppose we directly predict forces with error $e(x)=F_{\text{pred}}(x)-F(x)$. We can decompose the error in a conservative and a non-conservative field using the Helmholtz decomposition: $e(x)=\nabla\phi(x)+S(x)$. $\phi\in\mathbb{R}$ is a scalar potential, and S is a vector field $S(x)\in\mathbb{R}^{3N}$ that is divergence-free $\nabla\cdot S(x)=0$ but not curl-free $\exists x\text{ s.t. }\nabla\times S(x)\neq 0$. Now consider how the energy of the system changes over time in an MD simulation:
$\Delta E^{\prime}$ $=\int_{0}^{T}F_{\text{pred}}(x(t))v(t)dt=\int_{0}^{T}(F(x(t)+e(x(t)))v(t)dt$
$=\int_{0}^{T}(F(x)+\nabla\phi(x)+S(x))v(t)dt$
$=E(T)-E(0)+\underbrace{\phi(T)-\phi(0)}_{\text{ error bounded with time}}+\underbrace{\int_{0}^{T}S(x)v(t)dt}_{\text{error unbounded with time}}$
Since MD simulations use large $T$, non-conservative forces lead to the blowups in error.
In geometry optimization we define $E^{\prime}=E+\phi$, then follow the gradient descent ODE:
$\frac{dE^{\prime}(x(t))}{dt}$ $=\nabla E^{\prime}(x(t))\cdot\frac{dx(t)}{dt}$ (35)
$=-\nabla E^{\prime}(x(t))\cdot(\nabla E(x)+e(x))$ (36)
$=-\nabla E^{\prime}(x(t))\cdot(\nabla E(x)+\nabla\phi(x)+S(x))$ (37)
$=-\nabla E^{\prime}(x(t))\cdot(\nabla E^{\prime}(x)+S(x))$ (38)
$=-|\nabla E^{\prime}(x(t))|^{2}+\nabla E^{\prime}(x)\cdot S(x)$ (39)
Using the general inequality $2ab\leq a^{2}+b^{2}$:
$\frac{dE^{\prime}(x(t))}{dt}$ $=-|\nabla E^{\prime}(x(t))|^{2}+\nabla E^{\prime}(x)\cdot S(x)$ (40)
$\leq-|\nabla E^{\prime}(x(t))|^{2}+\frac{1}{2}|\nabla E^{\prime}(x(t))|^{2}+\frac{1}{2}|S(x)|^{2}$ (41)
$=-\frac{1}{2}|\nabla E^{\prime}(x(t))|^{2}+\frac{1}{2}|S(x)|^{2}$ (42)
Therefore, as long as $|\nabla E^{\prime}(x(t))|^{2}>|S(x)|^{2}$, the direct predicted force minimizes the energy plus some error potential similar to AD forces. Once we are close to convergence we have $|\nabla E^{\prime}(x(t))|^{2}\leq|S(x)|^{2}$ and stop converging or enter a limit cycle. This means, we will converge to an extrema on a energy surface $E^{\prime}$ within a distance of $O(|S|)$. In practice we are interested in how close we converge to the extrema of the DFT energy $E$. For this it doesnt matter if the error stems from the error potential $\phi$ shifting the extrema, or $S$ inducing a small limit cycle. As there is no time dependent error blow up it is therefore safe to use direct forces for geometry optimization.
We can perform the same analysis for the directly predicted Hessian $H_{\text{pred}}(x)=H(x)+e(x)$, and denote the error again as $e(x)=\nabla^{2}\phi(x)+S(x)$. As before we would get errors scaling with integration length if we would try to calculate quantities of the form
$Q(T)=\int_{0}^{T}f(x(t),H_{\text{pred}}(x(t)))dx(t)$ (43)
leading to large errors for large $T$. However, no mainstream task integrates dynamics over a function of the Hessian in such a manner. ZPE and geometry classification are single point calculation, therefore the errors $\nabla^{2}\phi(x)$ and $S(x)$ are both "constant".
Mode following in transition state search is equivalent to the geometry optimization analysis with direct forces above. Energy minimization with Hessian preconditioner is also unproblematic, as preconditioner cannot change the minimum, they only change the convergence speed. In fact, it is common practice to use BFGS updates for transition state search and energy minimization, and BFGS updates do not have any notion of integrability either. We conclude that it is save to use directly predicted Hessians for all mainstream applications.